Analysts expect MC, Co. to maintain a dividend payout ratio of 35% and enjoy an expected growth rate of 12% per year for the next 5 years. After the fifth year, all earnings will be paid out as dividends. The required rate of return on MC, Co equity is 8%.

At what price would the analysts value the stock under their own expectations?
Answer #1

Let us assume that face value per share is ₹100.

So dividend per share will be =100*35%=35(D0)

D1=D0(1+g)

=35(1+12/100)5

=61.68(value after 5 years)

P0=value of stock  

Formula =P0=D1/ke-g

61.68/0.08-0.12= -1542

Note- It is theoretical problem otherwise cost of equity cannot be less than growth rate.